OpenStudy (anonymous):

WILL MEDAL AND FAN Which ratios can form a proportion? A. 4/3, 3/4 B. 6/36, 5/30 C. 10/16, 16/24 D. 27/15, 11/31

OpenStudy (anonymous):

Can you multiply the numerator and denominator of the first number by the same number to get the second number? If you can, boom, proportion.

OpenStudy (anonymous):

Can you explain that more?

OpenStudy (anonymous):

So \[\large \sf \frac{a}{b} \times \frac{x}{x} = \frac{ax}{bx}\] Since the x's will cancel out, you are left with the original fraction a/b. Another way to think of it is "Can you simplify the second fraction to be the first one?" because if you can it is a ratio.

OpenStudy (anonymous):

Ok, thank you so much!!

OpenStudy (anonymous):

A third way would be if you have two fractions a/b and c/d, then \[\large \sf \frac{a}{b} = \frac{c}{d}\] means that when you cross multiply, they will be equal. \[\large \sf ad=bc\]

OpenStudy (anonymous):

Perfect thank you so so so much!!